If the sides of a square are lengthened by 7cm, the area becomes 121 cm squared. Find the length of a side of the original square.
I have no idea what I am doing, please help

Answer by ankor@dixie-net.com(22740)

If the sides of a square are lengthened by 7cm, the area becomes 121 cm squared.
Find the length of a side of the original square.
:
let x = side of the original square
then
(x+7) = side of the larger square which = 121 sq/cm
we can write this like this:
(x+7)^2 = 121
FOIL
x^2 + 14x + 49 = 121
:
x^2 + 14x + 49 - 121 = 0
:
x^2 + 14x - 72 = 0
Factor
(x + 18)(x - 4) = 0
Positive solution
x = 4 cm, size of the original square
:
:
Check: (4+7)^2 = 121
:
did this make sense to you now?